OpenStudy (anonymous):

0.26(x+40)+0.37(x+20)=-20

OpenStudy (tkhunny):

Were you planning to solve that? Let's see it!

OpenStudy (anonymous):

@tkhunny i dont know how to. i got the wrong answer i just want to see the steps. can you help me?

OpenStudy (tkhunny):

Doesn't work that way. Prove that you can do SOEMTHING. 0.26(x+40)+0.37(x+20)=-20 Do one piece at a time. 0.26(x+40) Can you expand that?

OpenStudy (anonymous):

.26x+10.4

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Perfect. Now the next piece. 0.37(x+20)

OpenStudy (anonymous):

@tkhunny .37x + 7.4

OpenStudy (tkhunny):

There you have it. Which part of that could you not do previously? Now add them all up! 0.26x + 10.4 + 0.37x + 7.4

OpenStudy (anonymous):

@tkhunny 3+0.11x=-20 -3 -3 0.11x=-23 divide 0.11x x=-209

OpenStudy (tkhunny):

0.26x + 0.37x = 0.63x 10.4 + 7.4 = 17.8 Why are you feeling a need to subtract? Did you write the problem statement incorrectly?

OpenStudy (anonymous):

@tkhunny ohhhhh okay thank you i get it. can you help me get through other ones please?

OpenStudy (anonymous):

2/5x-1/3x=2 @tkhunny

OpenStudy (tkhunny):

You'll have to be more clear. is "2/5x" this \(\dfrac{2}{5x}\) or this\(\dfrac{2}{5}x\)? Use parentheses to be clear.

OpenStudy (anonymous):

\[\frac{ 2 }{ 5 }x\]

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Now we add fractions? Denominators are 5 and 3. Go!